I need help! We're working with sequencing; I'm supposed to figure out which number comes next. I have no clue:
4, 6, 9, 13.5,?
12, -4, 4/3, -4/9, ?

Answer by Photonjohn(42)

For the first series, divide the first number by 2 and add it to that number, and so on. 4 +4/2 = 6. 6 + 6/2 = 9. 9 + 9/2 = 13.5 13.5 + 13.5/2 = ?
for the next series,
divide each number by -3 as the next number.
12/(-3) = -4
-4/(-3) = 4/3
(4/3)/(-3) = -4/9
etc.